a. At a certain temperature, the solubility of potassium iodate, KIO3, is 36.8 g/L. Calculate its solubility product constant for this temperature.

b. At a certain temperature, the solubility of cadmium fluoride, CdF2, is 0.185 mol/L. Calculate its solubility product constant for this temperature.

c. At a certain temperature, the solubility product constant* of copper (II) iodate, Cu(IO3)2, is 4.08x10-8 M3. Calculate the solubility of this compound for this temperature.
Answer #1

a)

Molar mass of KIO3,

MM = 1*MM(K) + 1*MM(I) + 3*MM(O)

= 1*39.1 + 1*126.9 + 3*16.0

= 214 g/mol

Molar mass of KIO3= 214 g/mol

s = 36.8 g/L

To covert it to mol/L, divide it by molar mass

s = 36.8 g/L / 214 g/mol

s = 0.172 mol/L

At equilibrium:

KIO3 <----> K+ + IO3-

   s s

Ksp = [K+][IO3-]

Ksp = (s)*(s)

Ksp = 1(s)^2

Ksp = 1(0.172)^2

Ksp = 2.957*10^-2

Answer: 2.96*10^-2

b)

At equilibrium:

CdF2 <----> Cd2+ + 2 F-

   s 2s

Ksp = [Cd2+][F-]^2

Ksp = (s)*(2s)^2

Ksp = 4(s)^3

Ksp = 4(0.185)^3

Ksp = 2.533*10^-2

Answer: 2.53*10^-2

C)

At equilibrium:

Cu(IO3)2 <----> Cu2+ + 2 IO3-

   s 2s

Ksp = [Cu2+][IO3-]^2

4.08*10^-8=(s)*(2s)^2

4.08*10^-8= 4(s)^3

s = 2.169*10^-3 M

Answer: 2.17*10^-3 M